OpenStudy (hba):

Find all the local maxima, local minima and saddle points of the function. f(x,y) = x e^y

OpenStudy (hba):

fx(x,y)=e^y fy(x,y)=xe^y e^y=0 x.e^y=0 y=1 x=0 (0,1) fxx(x,y)=0,fxy(x,y)=xe^y,fxy(x,y)=e^y D=-e^2 <0 Therfore,Saddle point.

OpenStudy (hba):

Please check

OpenStudy (yanasidlinskiy):

Wouldn't the question be something like this? f(x,y) = xe^(-2x^2 - 2y^2)

OpenStudy (math&ing001):

Both correct ! (0,1) is a local minima. D = fxx fyy - fxy^2 = -e^2 < 0 so (0,1) is a saddle point.

OpenStudy (hba):

How is it a local minima @math&ing001

OpenStudy (math&ing001):

You already found that it's a local minima/maxima. To check which one it is, plug in any random pair and compare it with f(0,1).